Choose the correct equation below which is set up such that the values of​ a, b, and c can be determined immediately.
A.
x^2+x=30
B.
(5x−3)(x−2)=0
C.
(5x+6)^2=7
D.
5x^2−32x−21=0
Solve the equation found in the previous step. The solution set is.
Please give a straight answer thank you.

Answer by stanbon(75887)

Choose the correct equation below which is set up such that the values of​ a, b, and c can be determined immediately.
A.
x^2+x=30
B.
(5x−3)(x−2)=0
C.
(5x+6)^2=7
D.
5x^2−32x−21=0
a = 5 ; b = -32 ; c = -21
-------------
Solve the equation found in the previous step. The solution set is.
Factor:
(x-7)(5x+3) = 0
x = 7 or x = -3/5
-------------
Cheers,
Stan H.
------------